### v1

# Decoherence Implies Information Gain

Abstract
It is shown that if the wave function of a quantum system undergoes an arbitrary random transformation such that the diagonal elements of the density matrix in the basis of a preferred observable remain constant, then (i) the off-diagonal elements of the density matrix become smaller in magnitude, and (ii) the system gains information about the preferred observable in the sense that the uncertainty of the observable reduces. These results are independent of the details of how the system-environment interaction generates the transformation of the state of the system, and together imply that decoherence leads to information gain, not information loss.
The phenomenon of decoherence – the reduction of the off-diagonal elements of the density matrix of a quantum system in the basis of some preferred observable – is undoubtedly one of the most important conceptual developments in the foundations of quantum mechanics over the past forty years. Although the debate continues over whether decoherence in itself resolves the measurement problem Adler2003 ; Ballentine , it does clearly provide deep insights into the nature of the transformations of the state of a quantum system that can occur when the system interacts with its environment JZ ; MS ; Zurek22 . It is widely held that decoherence represents the loss of information from the system into its environment Zurek1991 ; Zurek2009 . Here we challenge this view and show, in fact, that decoherence is about information flowing into the system from the environment, not the other way around. Our results are general, applying independently of the precise nature of system-environment interactions.
The present note will be organized as follows. First, we examine the traditional view that decoherence means information loss. Although this conclusion is based on what may appear to be a compelling line of reasoning, we point out that there are interpretational issues in the traditional argument. We show that if a quantum system gains information about a preferred observable, then the density matrix of the system necessarily decoheres in the basis of that observable. This is followed by a more general and perhaps surprising result, namely, that any random perturbation of the wave function that preserves the diagonal elements of the density matrix in the basis of a given observable necessarily gives rise to decoherence and that this necessarily implies information gain about that observable.
Let us begin by summarizing the argument leading to the “information-loss” interpretation of decoherence. The notion that all physical systems are ultimately open systems, facing an environment with which they interact, can be regarded as a fundamental principle of nature JZ1985 ; BP . With this in mind, we consider a quantum system that is initially in a pure state $|\psi_{0}\rangle$ that can be written in the form
$|\psi_{0}\rangle=\sum_{k=1}^{n}\sqrt{p_{k}}\,\mbox{$\rm e$}^{{\rm i}\theta_{k}}\,|X_{k}\rangle$ (1)
when expanded in terms of a set of normalized eigenstates $\{|X_{k}\rangle\}_{k=1,\,.\,.\,.\,,\,n}$ of a nondegenerate observable ${\hat{X}}$.
We are concerned in what follows with the overall magnitudes of the matrix elements of the initial density matrix
${\hat{r}}=|\psi_{0}\rangle\langle\psi_{0}|.$ (2)
The matrix elements of ${\hat{r}}$ in the $\hat{X}$-basis are given by
$r_{ij}=\langle X_{i}|\psi_{0}\rangle\langle\psi_{0}|X_{j}\rangle=\sqrt{p_{i}p_{j}}\,\mbox{$\rm e$}^{{\rm i}(\theta_{i}-\theta_{j})}$ (3)
for $i,j=1,\ldots,n$, whereas for the overall magnitudes one has
$|\,r_{ij}\,|=\left|\langle X_{i}|\psi_{0}\rangle\langle\psi_{0}|X_{j}\rangle\right|=\sqrt{p_{i}p_{j}}\,.$ (4)
In addition to the system, we consider an environment surrounding the system. The system then interacts with the environment, which holds many more degrees of freedom than the system. The system-environment interaction will then perturb the system to generate a transformation of the form
$|\psi_{0}\rangle\Rightarrow|\Psi\rangle=\sum_{k=1}^{n}\sqrt{\pi_{k}}\,\mbox{$\rm e$}^{{\rm i}\phi_{k}}|X_{k}\rangle.$ (5)
Now, an external observer typically has no access to detailed information about the perturbation. Hence, such an observer can at best merely speak about the statistics of the perturbed state. Therefore, we model the new probabilities $\{\pi_{i}\}$ and the new phases $\{\phi_{i}\}$ by taking them to be random variables. Since the expectation value of any observable ${\hat{F}}$ in a random state $|\Psi\rangle$ takes the form ${\rm tr}\,({\hat{F}}\,|\Psi\rangle\langle\Psi|)$, it follows that the state of the system, as far as any external observer is concerned, is given by the overall mean density matrix
${\hat{\rho}}={\mathbb{E}}\,[\,|\Psi\rangle\langle\Psi|\,],$ (6)
where ${\mathbb{E}}[-]$ denotes ensemble average over the random degrees of freedom. Thus, $|\Psi\rangle$ represents an “unravelling” of $\hat{\rho}$ Carmichael ; WD . We are interested in determining conditions on the random probabilities and phases in the transformation (5) sufficient to ensure that density matrix (6) exhibits decoherence. The matrix elements of ${\hat{\rho}}$, in the ${\hat{X}}$-basis, are
$\rho_{ij}={\mathbb{E}}\left[\langle X_{i}|\Psi\rangle\langle\Psi|X_{j}\rangle\right]={\mathbb{E}}\left[\!\sqrt{\pi_{i}\,\pi_{j}}\,\mbox{$\rm e$}^{{\rm i}(\phi_{i}-\phi_{j})}\right],$ (7)
for which the corresponding magnitudes $|\,\rho_{ij}\,|$ cannot in general be simplified, in the way we saw with (4), on account of the possible correlations between the random probabilities and the random phases. Indeed, the transformed probabilities $\{\pi_{i}\}_{i=1,\,.\,.\,.\,,\,n}$ and phases $\{\phi_{i}\}_{i=1,\,.\,.\,.\,,\,n}$ can take a variety of different forms, depending on the precise nature of the system-environment interaction.
For instance, in the special case of an environment that consists solely of an apparatus that detects the value of the observable ${\hat{X}}$, without recording the outcomes, the transformed probabilities take the form $\pi_{i}=\mathds{1}\{X=x_{i}\}$, where $X$ denotes the random variable corresponding to the outcome of the measurement. Here, for each value of $i$ we write ${\mathds{1}}\{X=x_{i}\}$ for the indicator function that takes the value one if $X=x_{i}$ and zero otherwise. The random phases are equal to their initial constant values in this example and for the random state vector representing the system after its interaction with the environment we obtain
$|\Psi\rangle=\sum_{k=1}^{n}{\mathds{1}}\{X=x_{k}\}\,\mbox{$\rm e$}^{{\rm i}\theta_{k}}|X_{k}\rangle.$ (8)
The initial density matrix of the system then has the matrix elements (3) in the ${\hat{X}}$-basis, whereas after the measurement the density matrix is diagonal in the ${\hat{X}}$-basis and takes the form
$\hat{\rho}=\sum_{k=1}^{n}p_{k}|X_{k}\rangle\langle X_{k}|.$ (9)
Thus the von Neumann entropy of the system increases in this example from the initial value of $-{\rm tr}\,(\hat{r}\log\hat{r})=0$ to a final value of $-{\rm tr}\,(\hat{\rho}\log\hat{\rho})=-\sum_{k}p_{k}\log p_{k}$ Clearly, there are countless other transformations that can be generated by a system-environment interaction. Whatever form these take, for decoherence the observable ${\hat{X}}$ plays an important role in this interaction. In other examples, the amount of increase of the von Neumann entropy will be different, but it remains true that the entropy increases, because the state (6) after the transformation is mixed, whereas the initial state is pure.
The argument then follows that the system must have lost information since the von Neumann entropy of the state of the system has increased. The environment, on the other hand, has gained this information, for instance, by detection of the measurement outcome – or so it is argued. That concludes the train of thought leading to the often-made claim that decoherence implies information loss. Let us point to some of the difficulties associated with this interpretation.
To begin, we recall that the notion of decoherence is basis dependent, whereas the von Neumann entropy is basis independent. Hence while decoherence is observed in the ${\hat{X}}$-basis, the information that is “lost” has no obvious relation to the preferred basis. This raises the question of exactly what information is lost by the system to the environment. Is it information about the state of the system? If so, this leads to the following issue. Namely, even if the environment has performed a measurement of ${\hat{X}}$, little information is obtained about the initial state of the system from the outcome. Some information is gained, for instance that the initial state $|\psi_{0}\rangle$ could not have been orthogonal to the output state, but mostly the information gained concerns the state of the system after the measurement.
With these considerations in mind, we offer an alternative interpretation of decoherence, based on the following points. First, we note that the “lost” information as measured by the increased von Neumann entropy is associated to an external observer, not the system. Initially the external observer is aware of the state of the system, but owing to random interactions between the system and its environment, the information held by an external observer is replaced by a statistical description of the state of the system, represented by a density matrix. In other words, the observer loses track of the state of the system. Further, the information about ${\hat{X}}$ is encoded in the Shannon-Wiener entropy associated with the squared magnitudes of the expansion coefficients of the state in the ${\hat{X}}$-basis, and it is this entropy that is more relevant to decoherence. We shall show that the difference between the initial Shannon-Wiener entropy and the averaged final Shannon-Wiener entropy is positive. Thus, the process of decoherence implies that the system gains information, not the environment.
We support our conclusions by proving four propositions. Proposition 1 asserts that if a quantum system gains information about an observable ${\hat{X}}$, then this leads to decoherence in the basis of ${\hat{X}}$. Proposition 2 is a more general result, namely, if a random transformation (5) preserves the diagonal elements of the density matrix in the basis of ${\hat{X}}$, then this leads to decoherence in that basis. This is perhaps surprising – or at least unexpected – since conventionally one associates decoherence with the decay of the off-diagonal elements; whereas here we show that there is a more fundamental characterization of decoherence purely in terms of the diagonal elements of the density matrix. In other words, there is no specific need to examine the off-diagonal elements of the density matrix to determine whether the system has decohered. Proposition 3, which arises as a corollary to Proposition 2, shows that when there is decoherence, the variance of ${\hat{X}}$ on average decreases, reflecting the fact that decoherence implies information gain by the system. Finally, in Proposition 4 we show that the Shannon-Wiener entropy decreases as a consequence of the interaction with the environment that generates decoherence.
To avoid some trivial cases, we shall assume that the random variables $\{\pi_{i}\}$ are linearly independent, that is to say, that no set of constants $\{\lambda_{i}\}$ exists such that
$\sum_{k=1}^{n}\lambda_{k}\pi_{k}=0.$ (10)
For instance, if $n=2$ then $\pi_{1}$ and $\pi_{2}$ are linearly dependent if and only if they are constant. This follows because there are two constraints $\pi_{1}+\pi_{2}=1$ and $\lambda_{1}\pi_{1}+\lambda_{2}\pi_{2}=0$ for some $\lambda_{1},\lambda_{2}$, so $\pi_{1}=\lambda_{2}/(\lambda_{2}-\lambda_{1})$ and $\pi_{2}=\lambda_{1}/(\lambda_{1}-\lambda_{2})$.
Proposition 1. Suppose that the the state of the system undergoes a random transformation of the form (5) in such a way that information is gained about the observable ${\hat{X}}$. Then the state of the system decoheres in the ${\hat{X}}$-basis.
Proof. We begin with a few preliminaries. For $p\geq 1$ we say as usual that a random variable $X$ lies in ${\cal L}^{p}$ if ${\mathbb{E}}[|X|^{p}]<\infty$. In this case its ${\cal L}^{p}$-norm is defined by $\|X\|_{p}=({\mathbb{E}}[|X|^{p}])^{1/p}$. Then we have the Schwarz inequality Williams , which in a probabilistic setting asserts that if $X,Y\in{\cal L}^{2}$ then $XY\in{\cal L}^{1}$ and
$\left|\,{\mathbb{E}}[XY]\,\right|\leq{\mathbb{E}}\left[\,|XY|\,\right]\leq\|X\|_{2}\,\|Y\|_{2}\,,$ (11)
with equality on the right if and only if $X$ and $Y$ are linearly dependent. A rather natural way of characterizing the notion of information gain in many settings is by use of Kolmogorov’s theory of conditional probabilities Williams . This theory has the advantage of enabling one to work at a high level of generality. Specifically, let ${\mathscr{I}}$ denote any choice of an information set (viz., a sigma-algebra) that is not independent of the random variable $X$. What non-independence means is that the specification of ${\mathscr{I}}$ will provide nontrivial information about the value of $X$. The corresponding informationally-rich random probabilities are given in terms of conditional expectations by
$\pi_{i}={\mathbb{P}}(X=x_{i}\,|{\mathscr{I}})={\mathbb{E}}[{\mathds{1}}\{X=x_{i}\}\,|{\mathscr{I}}].$ (12)
Then by the so-called tower property of conditional expectation we have
${\mathbb{E}}[\pi_{i}]={\mathbb{E}}[{\mathbb{E}}[{\mathds{1}}\{X=x_{i}\}\,|{\mathscr{I}}]]={\mathbb{E}}[{\mathds{1}}\{X=x_{i}\}]=p_{i}.$ (13)
Since the $\{\pi_{i}\}$ that are constructed in this way are centered at $\{p_{i}\}$ and linearly independent, and are such that $\sqrt{\pi_{i}}\in{\cal L}^{2}$ for each $i$, it follows that $\|\!\sqrt{\pi_{i}}\,\|_{2}=\sqrt{p_{i}}$, and from (11) we have
${\mathbb{E}}\left[\!\sqrt{\pi_{i}\pi_{j}}\,\right]\,<\,\|\!\sqrt{\pi_{i}}\,\|_{2}\,\,\|\!\sqrt{\pi_{j}}\,\|_{2}=\!\sqrt{p_{i}p_{j}}\,,$ (14)
for $i\neq j$. Next, we consider the magnitudes of the matrix elements (7) of the transformed state, given by
$|\,\rho_{ij}\,|=\left|\,{\mathbb{E}}\left[\!\sqrt{\pi_{i}\,\pi_{j}}\,\mbox{$\rm e$}^{{\rm i}(\phi_{i}-\phi_{j})}\right]\,\right|.$ (15)
We remark that if $Z$ is any complex-valued random variable with integrable real and imaginary parts, then
${\mathbb{E}}[\,|\,Z\,|\,]\geq|\,{\mathbb{E}}[Z]\,|,$ (16)
with equality only if $Z$ is constant. This follows by use of Jensen’s inequality for a convex function of two real variables. Since the argument of the expectation in (15) is a complex random variable, we can use (16) to deduce that
$|\,\rho_{ij}\,|\leq\,{\mathbb{E}}\left[\,\left|\!\sqrt{\pi_{i}\,\pi_{j}}\,\mbox{$\rm e$}^{{\rm i}(\phi_{i}-\phi_{j})}\right|\,\right]={\mathbb{E}}\left[\!\sqrt{\pi_{i}\pi_{j}}\,\right],$ (17)
for $i\neq j$. Finally, if we combine (4), (14), and (17), we get
$|\,\rho_{ij}\,|\leq{\mathbb{E}}\left[\!\sqrt{\pi_{i}\pi_{j}}\,\right]<\!\sqrt{p_{i}p_{j}}=|\,r_{ij}\,|\,,$ (18)
for $i\neq j$ , which shows that the magnitudes of the off-diagonal matrix elements of the state after interaction with the environment will be strictly less than the corresponding magnitudes for the initial state, as claimed. $\Box$
It is interesting to note that if the random phases are statistically independent from the random probabilities, then the inequality on the left side of (17) becomes an equality. Nonetheless, we still obtain (18) for $i\!\neq\!j$. In other words, if the amplitudes and phases are statistically dependent, then the inequalities become even sharper. A special case of Proposition 1 was established in BT in the situation where the information set ${\mathscr{I}}$ is generated by a random variable that is correlated to $X$. Here we have generalized that result to the case where the conditioning is with respect to an arbitrary sigma-algebra.
Proposition 2. Suppose that the quantum state undergoes a transformation of the form (5) such that the probabilities are conserved in the basis of an observable ${\hat{X}}$, so that
${\mathbb{E}}[\pi_{i}]=p_{i}\,.$ (19)
Then we obtain (18) for all $i\neq j$.
Proof. Clearly $\sqrt{\pi_{i}}\in{\cal L}^{2}$. Further, by use of (19) we have $\|\!\sqrt{\pi_{i}}\,\|_{2}=\sqrt{p_{i}}$. Hence, by the Schwarz inequality (11) and the complex inequality (16) we get (18), as claimed. $\Box$
Though Propositions 1 and 2 appear similar, there is a subtle difference. In Proposition 1, we assume that the gain in information is achieved by conditioning with respect to an information set, which automatically implies (19). In Proposition 2 we have not stipulated a particular mechanism for the information gain – we have merely assumed the mean condition. Therefore, if a quantum state undergoes a random transformation such that the diagonal elements of the density matrix are conserved in expectation when expressed in the basis of ${\hat{X}}$, then the state will decohere in that basis.
Proposition 3. Let the quantum state undergo a transformation (5) such that (19) holds. Then
${\mathbb{E}}\left[\langle\Psi|{\hat{X}}^{2}|\Psi\rangle-\langle\Psi|{\hat{X}}|\Psi\rangle^{2}\right]<\langle\psi_{0}|{\hat{X}}^{2}|\psi_{0}\rangle-\langle\psi_{0}|{\hat{X}}|\psi_{0}\rangle^{2},$ (20)
showing that the averaged uncertainty of ${\hat{X}}$ is reduced.
Proof. For $W\in{\cal L}^{2}$ we have ${\mathbb{E}}[W^{2}]\geq\left({\mathbb{E}}[W]\right)^{2}$ by Jensen’s inequality, with equality if and only if $W$ is constant. Hence if we set $W=\sum_{k}\pi_{k}x_{k}$, which is square-integrable, we obtain
${\mathbb{E}}\left[\,\left(\sum_{k=1}^{n}\pi_{k}x_{k}\right)^{2}\right]>\left({\mathbb{E}}\left[\sum_{k=1}^{n}\pi_{k}x_{k}\right]\,\right)^{2}=\left(\sum_{k=1}^{n}p_{k}x_{k}\right)^{2}\!,$ (21)
by use of the mean condition (19) and our assumption that the $\{\pi_{i}\}$ are linearly independent. Changing the sign and adding equal terms to each side, we obtain
${\mathbb{E}}\left[\sum_{k=1}^{n}\pi_{k}x_{k}^{2}\right]-{\mathbb{E}}\left[\left(\sum_{k=1}^{n}\pi_{k}x_{k}\right)^{2}\right]<\sum_{k=1}^{n}p_{k}x_{k}^{2}-\left(\sum_{k=1}^{n}p_{k}x_{k}\right)^{2}\!,$ (22)
again by use of (19). But this gives (20), as claimed. $\Box$
The fact that the uncertainty on average reduces is another way of illustrating information gain. In BT ; DCB it was conjectured that if a quantum system decoheres in the basis of an observable ${\hat{X}}$, then the variance of ${\hat{X}}$ will on average decrease. This conjecture was shown in BT to be true in the case of a two-level system, whereas here we have generalized the result to systems of any size, with a considerably simpler proof.
Another way of characterizing the uncertainty of ${\hat{X}}$ in the state is in terms of the Shannon-Wiener entropy. The initial value of the Shannon-Wiener entropy of the system is
$S_{0}=-\sum_{k}p_{k}\log p_{k}\,,$ (23)
whereas for the final value on average we have
$S_{\!f}=-{\mathbb{E}}\,\left[\sum_{k}\pi_{k}\log\pi_{k}\right].$ (24)
Proposition 4. Suppose that a quantum state undergoes a transformation of the form (5) such that (19) holds. Then
$S_{\!f}<S_{\!0}\,.$ (25)
Proof. It is not difficult to check that the function defined by $I(x)=x\log x$ for $x>0$ is convex. Therefore, for any strictly positive integrable random variable $W$ it holds that
${\mathbb{E}}[W\log W]\geq{\mathbb{E}}[W]\log{\mathbb{E}}[W]$ (26)
by Jensen’s inequality, with equality only if $W$ is constant. It follows by use of (19) that
${\mathbb{E}}[\pi_{k}\log\pi_{k}]>{\mathbb{E}}[\pi_{k}]\log{\mathbb{E}}[\pi_{k}]=p_{k}\log p_{k},$ (27)
for each $k=1,.\,.\,.\,,n$, and hence
$-{\mathbb{E}}\,\left[\sum_{k}\pi_{k}\log\pi_{k}\right]<-\sum_{k}p_{k}\log p_{k},$ (28)
as claimed. Therefore, under any random transformation of the state that preserves the diagonal elements of the density matrix in the basis of a preferred observable ${\hat{X}}$, the averaged entropy of the state for the observable ${\hat{X}}$ decreases from its initial value following an interaction with the environment. $\Box$
Our results show that we can model the decoherence of a quantum system due to interactions with the environment by a randomization of the state of the system that in expectation conserves the probability law for the outcome of the measurement of a preferred observable. In fact, one can take this property of conservation of probability to be a definition of what we mean by an interaction with the environment that generates decoherence. Then decoherence in the basis of the preferred observable implies that the system has on average acquired information about that observable, as quantified by a decrease in the Shannon-Wiener entropy associated with the squared expansion coefficients of the state in that basis.
In modelling the state transition by a random perturbation, we step outside of the usual formalism of quantum mechanics with the inclusion of random elements. This can be justified on the basis that we are in essence putting forward a reduced-form model for decoherence, rather than, say, proposing a fundamentally new dynamics for the state vector. One can take a similar view on the dynamics of collapse models, since these also show decoherence of the density matrix while preserving the mean condition ABBH2001 ; BH2002 ; BH2023 .
In summary, in the present work we have shown that decoherence implies information gain by the quantum system, rather than information loss, contrary to conventional wisdom, and that the decay of the off-diagonal elements of the density matrix follows from a deep conservation law that preserves the diagonal elements of the density matrix. We hope that our findings will lead to new ways of understanding the emergence of classicality from an informational perspective.
Acknowledgements. The authors thank Simon Saunders for stimulating discussions. DCB acknowledges support from EPSRC (EP/X019926) and the John Templeton Foundation (grant 62210). The views expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation.